OpenStudy (anonymous):

g(x)=4x^2-4 is translated vertically downward by 5 units, it becomes the graph of a function F Find the expression for f(x)

jimthompson5910 (jim_thompson5910):

since g(x) "is translated vertically downward by 5 units", this means we're subtracting 5 from g(x) to get g(x)-5 But g(x) = 4x^2-4 So g(x)-5 = (4x^2-4)-5 = 4x^2-9 This means that g(x)-5 = 4x^2-9 and f(x) = 4x^2-9